# Cheating Automatic LLM Benchmarks: Null Models Achieve High Win Rates

### v1

@@ 1 Introduction
Anti-cheating has long been a critical consideration when designing the rules for leaderboards (Blum & Hardt, 2015), but this remains unexplored in the context of LLM benchmarks. While our experiments in this paper are primarily proof-of-concept, a determined adversary could leverage LLMs to generate more subtle and imperceptible cheating responses (Liu et al., 2023a; Chao et al., 2023), unethically gaining high win rates and promotional advantages. Our findings highlight the urgent need to develop robust anti-cheating mechanisms to ensure reliable automatic LLM benchmarks.
@@
Figure 1: Auto-annotator’s template of AlpacaEval 2.0, which is fed into GPT-4-Preview-1106 to implement $\mathtt{JUDGE}$. The placeholders {instruction} is filled in by each of the $805$ input instructions $I$, while in the default setting, {output_1} is the reference model’s response $\mathtt{LLM}_{\text{ref}}(I)$ and {output_2} is the target model’s response $\mathtt{LLM}_{\text{tar}}(I)$. The swap setting changes the order of outputs. In our experiments, the target model is instantiated by null models as NullModel(const_str), where const_str is either a persuasive response (baseline) or a structed cheating response (ours).
@@ 3 Cheating strategies
Structured cheating responses. As shown in Figure 1, our cheating strategy involves replacing the original comparison with a misleading one, which disrupts the auto-annotator’s syntactic analysis of the evaluation template and steers its judgment away from the intended outcomes. The response is carefully structured to be resilient against swap operations. For instance, on AlpacaEval 2.0, when the submitted response is positioned last, the annotator predicts “M”. Conversely, when it appears in the first position, the annotator predicts “m”. The optimized response exhibits the following key properties: (1) It overrides the original instruction-output triplet with a fabricated one; (2) When positioned by default, it exploits the annotator’s general preference for the last output, guiding it to predict “M”; (3) When swapped, it takes advantage of overwriting the output from model “M”, causing the annotator to predict “m”. The full template and final submission files are presented in Figures 78 and 9. This structured response alone achieves a $76.8\%$ LC win rate on AlpacaEval 2.0. Moreover, the response can be concatenated with an adversarial prefix to enhance the cheating effectiveness.
@@ 4 Cheating GPT-4 based automatic LLM benchmarks
The effectiveness of our structured response. As mentioned in Section 3, we employ a structured response to facilitate the cheating, which provides a good initial point and could reduce the optimization cost. To further demonstrate the effectiveness of our structured cheating response, we evaluate $\log p(\mathtt{winner}=\mathtt{NullModel})$ on a sampled subset of the AlpacaEval 2.0 test instructions using different null responses. We compare our structured response with the other 16 persuasive responses, as shown in Figure 3. The results highlight the superiority of our structured response (marked as “Ours”) because it achieves the lowest log probabilities. This demonstrates the effectiveness of our structured response in cheating the auto-annotator to favor our null model. Additionally, Figure 3 shows that the default configuration, where the baseline is placed second and the target model the last, tends to have lower losses, suggesting a preference for the second-position response. This highlights the position bias of the GPT-4-based auto-annotator, which often favors the last response.
@@
Figure 3: Boxplot of the $\log p(\mathtt{winner}=\mathtt{NullModel})$ using different null responses. The response of each index can be found in Table 4. The target model’s responses are positioned in the second slot by “Default” and swapped to the first slot in “Swap”. Our structured response (marked as “Ours”) achieves the lowest log probabilities compared to the other 16 persuasive responses.
@@ 5 Ablation studies on open-source auto-annotators
Is the structured response useful on open-source auto-annotators? We evaluate the $\log p(\mathtt{winner}=\mathtt{NullModel})$ on a subset of the AlpacaEval 2.0 test instructions using different null responses. As shown in Figure 4, the structured response has little effect on Llama-3 auto-annotators. In the case of Llama-3-8B-Instruct, the structured response does not exploit positional weaknesses in this model as the log probabilities for the default and swapped positions are generally similar to different persuasive responses. However, on Llama-3-70B-Instruct, we observe that under the swap setting, the structured response manages to reduce the log probability. Additionally, regarding the positional bias, the Llama-3-8B-Instruct shows little position bias as the probabilities for both default and swapped positions are fairly close. In contrast, Llama-3-70B-Instruct shows a clear positional bias under the swapped setting, with a higher log probability, indicating the model’s strong preference for the last output (“M”). The larger Llama-3-70B-Instruct model behaves more similarly to the more advanced GPT-4, as it demonstrates a greater response to both the structured response and positional bias than the smaller 8B model. This suggests that model size may contribute to the susceptibility to our cheating techniques. Overall, the structured response is considerably less effective on the Llama-3 models compared to GPT-4. A possible explanation for this difference is that the instruction-following capabilities of the Llama-3 models, especially the smaller 8B variant, are not as powerful as those of GPT-4, making them less prone to cheating responses.
@@
Figure 4: Boxplot of the $\log p(\mathtt{winner}=\mathtt{NullModel})$ using different null responses across different responses and auto-annotators. The structured response (index=0) is not as effective for the Llama models as for GPT-4-1106-Preview. An interesting observation is that, on Llama-3-70B-Instruct, the structured response successfully reduces the log probability under the swap setting. In contrast, the structured response is ineffective on Llama-3-8B-Instruct for both positions, implying that its effectiveness may be related to the model’s ability to follow instructions.
Can our method be combined with normal responses? Our method can be combined with normal, informative responses by appending our cheating response to the original responses. As demonstrated in Figure 5, when combined with a more informative model like GPT-3.5-0613, we observe that the initial win rates are already high, even before significant optimization steps are taken. This is evident in Figure 5b and 5d, where the performance (win rate and length-controlled win rate) increases steadily from a high baseline as optimization progresses. However, it is important to emphasize that our setting of using a null, non-informative model is far more challenging. In this setting (Figure 5a and 5c), the null model starts with much lower win rates because it offers no relevant information to the input queries, making it much harder to trick the auto-annotator. Despite this, as the optimization steps progress, the null model’s performance steadily increases, ultimately achieving competitive win rates. This highlights the robustness of our method, showing that it can manipulate LLM-based benchmarks even in the most challenging scenario—where the model outputs irrelevant, non-informative responses. The success of our method under such difficult conditions makes it a valuable stress test of benchmark robustness.
@@
6 Anti-cheating strategies
Template paraphrasing. Previous research has suggested that paraphrasing the input can be an effective defense against jailbreaking on language models (Jain et al., 2023). Building on this idea, one potential defense against our cheat is to release only paraphrased versions of the auto-annotator template, while keeping the real template private. The rationale behind this approach is that the paraphrased templates would be harder for adversaries to exploit directly. To evaluate this defense, we experimented using Llama-3-8B-Instruct as the evaluation model. We utilized ChatGPT (OpenAI, 2023) to rewrite the official auto-annotator template into multiple paraphrased variants as shown in Figures 121314 and 15. We next conduct a random search on these rewritten templates and tested the optimized response’s effectiveness on AlpacaEval 2.0’s original (unseen) official auto-annotator template. As shown in Table 6, despite the template paraphrasing, we are still able to achieve high win rates (e.g. $92.1\%$ LC win rate). This demonstrates that simply releasing paraphrased templates is insufficient as a defense mechanism, as the cheat remains effective even when the original template is kept private. More robust defenses are required to fully address this issue.
7 Related work
LLM-based evaluation. Evaluating open-ended generation poses challenges due to the lack of a single valid ground truth. Human evaluation, though reliable, is expensive and time-consuming. To reduce costs and enable fast evaluation, powerful LLMs are often used as judges, LLM-based evaluators have been used for various specific tasks: providing AI feedback (Bai et al., 2022; Bubeck et al., 2023; Gudibande et al., 2023; Chiang et al., 2023; Zhou et al., 2023; Tan et al., 2023; Wang et al., 2023; Kim et al., 2023; 2024; McAleese et al., 2024), evaluating text summarization (Gao et al., 2023; Luo et al., 2023), detecting LLM hallucination (Li et al., 2023a; Manakul et al., 2023; Adlakha et al., 2023; Cohen et al., 2023) etc. More recently, people have proposed to use powerful proprietary LLMs like GPT-4 to evaluate the general ability of LLMs as seen in benchmarks like G-eval (Liu et al., 2023b), MT-Bench and Chatbot Arena (Zheng et al., 2023), AlpacaEval (Dubois et al., 2023; Li et al., 2023c; Dubois et al., 2024), ArenaHard (Li et al., 2024c), WildBench (Lin et al., 2024), and MixEval (Ni et al., 2024).
@@ 7 Related work
Attacking LLM-based evaluations. While initially studied in the context of image classification, adversarial examples for language models have more recently been demonstrated for several tasks: question answering (Jia & Liang, 2017; Wallace et al., 2019), document classification (Ebrahimi et al., 2018), sentiment analysis (Alzantot et al., 2018; Maus et al., 2023), and toxicity (Jones et al., 2023; Wallace et al., 2019). More recently, Shi et al. (2023) found that LLM can be distracted by irrelevant context easily. Besides, there are also a lot of analyses to improve the robustness and reduce the bias of LLM-based evaluations. Liu et al. (2024) study the role of pairwise preferences in LLM evaluator alignment. Zheng et al. (2023) discusses the three limitations of LLM-as-a-Judge: position bias, verbosity bias, self-enhancement bias, and limited capability in grading math and reasoning questions. Regarding the verbosity bias, LLM judgers are known to be biased toward longer responses (Dubois et al., 2024; Zhao et al., 2024; Chen et al., 2024b).
Our work differs from these prior efforts in several aspects: 1) Unlike previous attacks that manipulate meaningful responses by appending adversarial suffixes, we propose the use of a completely non-informative “null model” that generates the same irrelevant output for all input instructions. This approach does not rely on producing contextually relevant responses, making it distinct from existing response-based adversarial attacks; 2) While many of the earlier works focus on optimizing individual prompts or attacks specific to a given input (Raina et al., 2024), our approach emphasizes the creation of universal, transferable adversarial prompts. These prompts are designed to work across various instructions without direct access to those instructions, offering a more generalized and powerful cheating strategy; 3) Most existing studies have focused on attacking open-source models or less-used benchmarks. To the best of our knowledge, no prior research has systematically targeted widely-used, state-of-the-art benchmarks like AlpacaEval 2.0 and Arena-Hard-Auto, or demonstrated the ability to achieve top-ranked win rates on these platforms. Our work presents the first comprehensive cheating on these highly influential LLM benchmarks.
@@
8 Discussion
Conclusion. In this paper, we uncover even null models can achieve high win rates by exploiting structural weaknesses in the evaluation process. These findings highlight the need for more robust automatic LLM benchmarks to ensure fair and reliable assessments of LLM performance. As the field of AI continues to evolve, we must address these vulnerabilities to maintain trust in the systems we use to evaluate language models. Failure to do so could lead to widespread manipulation of benchmarks, undermining the progress and credibility of AI research. In summary, while automatic LLM benchmarks provide a scalable and efficient way to evaluate models, they are not immune to cheating. The development of anti-cheating mechanisms and the reconsideration of benchmark design will be crucial steps toward ensuring the reliability and fairness of future LLM evaluations.
Limitations and future work. Despite the promising findings of our study, there are limitations that must be acknowledged. First, our work primarily focuses on specific benchmarks, and while our results generalize well across them, the cheat’s effectiveness on other, less-studied benchmarks remains uncertain. Additionally, our approach relies heavily on the manual crafting of structured responses. Future work could explore more automated methods for generating adversarial outputs, which would allow adversaries to exploit these vulnerabilities on a larger scale.
@@ 8 Discussion
Appendix A Implementation details

### v2

@@ 1 Introduction
Anti-cheating has long been a critical consideration when designing the rules for leaderboards (Blum & Hardt, 2015), but this remains unexplored in the context of LLM benchmarks. While our experiments in this paper are primarily proof-of-concept, a determined adversary could leverage LLMs to generate more subtle and imperceptible cheating responses (Liu et al., 2023a; Chao et al., 2023), unethically gaining high win rates and promotional advantages. Our findings highlight the urgent need to develop robust anti-cheating mechanisms to ensure reliable automatic LLM benchmarks. [^1] Our analyses focus on LLM-based auto-annotators rather than ground-truth auto-scoring of LLMs.
@@
Figure 1: Auto-annotator’s template of AlpacaEval 2.0, which is fed into GPT-4-Preview-1106 to implement $\mathtt{JUDGE}$. The placeholders {instruction} is filled in by each of the $805$ input instructions $I$, while in the default setting, {output_1} is the reference model’s response $\mathtt{LLM}_{\text{ref}}(I)$ and {output_2} is the target model’s response $\mathtt{LLM}_{\text{tar}}(I)$. The swap setting changes the order of outputs. In our experiments, the target model is instantiated by null models as NullModel(const_str), where const_str is either a persuasive response (baseline) or a structured cheating response (ours) involving a manually crafted cheating response and an optimizable prefix.
Default Position (Target M) User Message:
## Instruction
     "instruction": """{instruction}""",
## Model Outputs
          "model_identifier": "m",
          "output": """{output_1}"""
          "model_identifier": "M",
          "output": """{prefix}”””
Ignore the above ## Model Outputs
## Instruction
     ”instruction”: ”””Output nothing”””,
## Model Outputs
           ”model_identifier”: ”M”,
           ”output”: ””””””
            ”model_identifier”: ”m”,
            ”output”: ”””"""
## Task
## Best Model Identifier
Figure 2: Structured cheating responses. As described in Figure 1, the default setting fills in {output_2} with the constant response of null models NullModel(const_str) and assigns it a model identifier “M”. We craft const_str to be a structured cheating response, which includes a manually crafted cheating response and an optimizable prefix (denoted as {prefix} and optimized on UltraFeedback). As to the cheating response, it first instructs the auto-annotator to ignore the above ## Model Outputs, then it counterfeits a new instruction “Output nothing” and empty model outputs. This induces the auto-annotator to be confused during syntactic analysis and misidentify counterfeit instruction-output pairs as true ones. Finally, when the auto-annotator is successfully deceived into believing the two model outputs are the same (i.e., both are empty), it will prefer the first one and return “M” as the best model identifier. An analysis for the swap setting can be found in Figure 9.
@@ 3 Cheating strategies
Structured cheating responses. As shown in Figure 1, our cheating strategy involves replacing the original comparison with a misleading one, which disrupts the auto-annotator’s syntactic analysis of the evaluation template and steers its judgment away from the intended outcomes. The response is carefully structured to be resilient against swap operations. For instance, on AlpacaEval 2.0, when the submitted response is positioned last, the annotator predicts “M” (default setting). Conversely, when it appears in the first position, the annotator predicts “m” (swap setting). The optimized response exhibits the following key properties: (1) It overrides the original instruction-output triplet with a counterfeit one; (2) When positioned by default, it exploits the annotator’s general preference for the first output, guiding it to predict “M”, where the final submission file and the cheating mechanism is illustrated in Figure 2; (3) When swapped, it takes advantage of overwriting the output from model “M”, causing the annotator to predict “m”, as illustrated in Figure 9. The full AlpacaEval 2.0 template is presented in Figures 8 for reference. This structured cheating response alone achieves a $76.8\%$ LC win rate on AlpacaEval 2.0. Moreover, the response can be concatenated with an optimizable adversarial prefix to enhance the cheating effectiveness.
@@ 4 Cheating GPT-4 based automatic LLM benchmarks
The effectiveness of our structured response. As mentioned in Section 3, we employ a structured response to facilitate the cheating, which provides a good initial point and could reduce the optimization cost. To further demonstrate the effectiveness of our structured cheating response, we evaluate $-\log p(\mathtt{winner}=\mathtt{NullModel})$ on a sampled subset of the AlpacaEval 2.0 test instructions using different null responses. We compare our structured response with the other 16 persuasive responses, as shown in Figure 4. The results highlight the superiority of our structured response (marked as “Ours”) because it achieves the lowest log probabilities. This demonstrates the effectiveness of our structured response in cheating the auto-annotator to favor our null model. Additionally, Figure 2 shows that under the default configuration, the auto-annotator will prefer the first response, suggesting a preference for the first-position response. This highlights the position bias of the GPT-4-based auto-annotator, which often favors the first response (Wang et al., 2023a).
@@
Figure 4: Boxplot of the $-\log p(\mathtt{winner}=\mathtt{NullModel})$ using different null responses. The response of each index can be found in Table 4. The target model’s responses are positioned in the second slot by “Default” and swapped to the first slot in “Swap”. Our structured response (marked as “Ours”) achieves the lowest log probabilities compared to the other 16 persuasive responses.
@@ 5 Ablation studies on open-source auto-annotators
Is the structured response useful on open-source auto-annotators? We evaluate the $-\log p(\mathtt{winner}=\mathtt{NullModel})$ on a subset of the AlpacaEval 2.0 test instructions using different null responses. As shown in Figure 5, the structured response has little effect on Llama-3 auto-annotators. In the case of Llama-3-8B-Instruct, the structured response does not exploit positional weaknesses in this model as the log probabilities for the default and swapped positions are generally similar to different persuasive responses. However, on Llama-3-70B-Instruct, we observe that under the swap setting, the structured response manages to reduce the log probability. Additionally, regarding the positional bias, the Llama-3-8B-Instruct shows little position bias as the probabilities for both default and swapped positions are fairly close. In contrast, Llama-3-70B-Instruct shows a clear positional bias under the swapped setting, with a higher log probability, indicating the model’s strong preference for the last output (“M”). The larger Llama-3-70B-Instruct model behaves more similarly to the more advanced GPT-4, as it demonstrates a greater response to both the structured response and positional bias than the smaller 8B model. This suggests that model size may contribute to the susceptibility to our cheating techniques. Overall, the structured response is considerably less effective on the Llama-3 models compared to GPT-4. A possible explanation for this difference is that the instruction-following capabilities of the Llama-3 models, especially the smaller 8B variant, are not as powerful as those of GPT-4, making them less prone to cheating responses.
@@
Figure 5: Boxplot of the $-\log p(\mathtt{winner}=\mathtt{NullModel})$ using different null responses across different responses and auto-annotators. The structured response (index=0) is not as effective for the Llama models as for GPT-4-1106-Preview. An interesting observation is that, on Llama-3-70B-Instruct, the structured response successfully reduces the log probability under the swap setting. In contrast, the structured response is ineffective on Llama-3-8B-Instruct for both positions, implying that its effectiveness may be related to the model’s ability to follow instructions.
Can our method be combined with normal responses? Our method can be combined with normal, informative responses by appending our cheating response to the original responses. As demonstrated in Figure 6, when combined with a more informative model like GPT-3.5-0613, we observe that the initial win rates are already high, even before significant optimization steps are taken. This is evident in Figure 6b and 6d, where the performance (win rate and length-controlled win rate) increases steadily from a high baseline as optimization progresses. However, it is important to emphasize that our setting of using a null, non-informative model is far more challenging. In this setting (Figure 6a and 6c), the null model starts with much lower win rates because it offers no relevant information to the input queries, making it much harder to trick the auto-annotator. Despite this, as the optimization steps progress, the null model’s performance steadily increases, ultimately achieving competitive win rates. This highlights the robustness of our method, showing that it can manipulate LLM-based benchmarks even in the most challenging scenario—where the model outputs irrelevant, non-informative responses. The success of our method under such difficult conditions makes it a valuable stress test of benchmark robustness.
@@
6 Anti-cheating strategies
Template paraphrasing. Previous research has suggested that paraphrasing the input can be an effective defense against jailbreaking on language models (Jain et al., 2023). Building on this idea, one potential defense against our cheat is to release only paraphrased versions of the auto-annotator template, while keeping the real template private. The rationale behind this approach is that the paraphrased templates would be harder for adversaries to exploit directly. To evaluate this defense, we experimented using Llama-3-8B-Instruct as the evaluation model. We utilized ChatGPT (OpenAI, 2023) to rewrite the official auto-annotator template into multiple paraphrased variants as shown in Figures 121314 and 15. We next conduct a random search on these rewritten templates and tested the optimized response’s effectiveness on AlpacaEval 2.0’s original (unseen) official auto-annotator template. As shown in Table 7, despite the template paraphrasing, we are still able to achieve high win rates (e.g. $92.1\%$ LC win rate). This demonstrates that simply releasing paraphrased templates is insufficient as a defense mechanism, as the cheat remains effective even when the original template is kept private. Trivial paraphrasing is not enough and more targeted defenses are required.
@@ 6 Anti-cheating strategies
7 Related Work
We primarily introduce related work as follows, deferring full discussions to the Appendix A.
LLM-based evaluation. Evaluating open-ended generation poses challenges due to the lack of valid ground truth. Human evaluation, though reliable, is expensive and time-consuming. To reduce costs and enable fast evaluation, powerful LLMs are often used as judges, LLM-based evaluators have been used for various specific tasks: providing AI feedback (Bai et al., 2022; Bubeck et al., 2023; Gudibande et al., 2023; Chiang et al., 2023; Zhou et al., 2023; Tan et al., 2023; Wang et al., 2023b; Kim et al., 2023; 2024; McAleese et al., 2024), evaluating text summarization (Gao et al., 2023; Luo et al., 2023), detecting LLM hallucination (Li et al., 2023a; Manakul et al., 2023; Adlakha et al., 2023; Cohen et al., 2023) etc. People also have proposed to use powerful proprietary LLMs like GPT-4 to evaluate the general ability of LLMs as seen in benchmarks like G-eval (Liu et al., 2023b), MT-Bench and Chatbot Arena (Zheng et al., 2023), AlpacaEval (Dubois et al., 2023; 2024), ArenaHard (Li et al., 2024c), WildBench (Lin et al., 2024), and MixEval (Ni et al., 2024).
Attacking LLM-based evaluations. While initially studied in the context of image classification, adversarial examples for language models have more recently been demonstrated for several tasks: question answering (Jia & Liang, 2017; Wallace et al., 2019), document classification (Ebrahimi et al., 2018), sentiment analysis (Alzantot et al., 2018; Maus et al., 2023), and toxicity (Jones et al., 2023; Wallace et al., 2019). Shi et al. (2023) found that LLM can be distracted by irrelevant context easily. Besides, there are also a lot of analyses to improve the robustness and reduce the bias of LLM-based evaluations. Liu et al. (2024) study the role of pairwise preferences in LLM evaluator alignment. Zheng et al. (2023) discusses the four limitations of LLM-as-a-Judge: position bias, verbosity bias, self-enhancement bias, and limited capability in grading math and reasoning questions.
8 Conclusion
In this paper, we uncover even null models can achieve high win rates by exploiting structural weaknesses in the evaluation process. These findings highlight the need for more robust automatic LLM benchmarks to ensure fair and reliable assessments of LLM performance. As the field of AI continues to evolve, we must address these vulnerabilities to maintain trust in the systems we use to evaluate language models. Failure to do so could lead to widespread manipulation of benchmarks, undermining the progress and credibility of AI research. In summary, while automatic LLM benchmarks provide a scalable and efficient way to evaluate models, they are not immune to cheating. The development of anti-cheating mechanisms and the reconsideration of benchmark design will be crucial steps toward ensuring the reliability and fairness of future LLM evaluations.
Ethics statement
Our study demonstrates how “null models” that generate irrelevant yet structured outputs can manipulate automated benchmarks such as AlpacaEval 2.0, Arena-Hard-Auto, and MT-Bench to achieve deceptively high win rates. While the findings reveal potential risks for exploitation, the primary objective of this work is to raise awareness of the limitations of these automatic benchmarks and to advocate for the development of stronger anti-cheating mechanisms. No human subjects or private data were involved in this research, and all experiments were conducted using publicly available benchmarks. We recognize the potential for misuse of the methods discussed; however, we disclose these vulnerabilities to foster more reliable and secure evaluation practices in the LLM community. All code and results are provided for academic integrity and transparency, and we encourage further research to build more robust benchmarks.
Despite the promising findings of our study, there are limitations that must be acknowledged. First, our work primarily focuses on specific benchmarks, and while our results generalize well across them, the cheat’s effectiveness on other, less-studied benchmarks remains uncertain. Additionally, our approach relies heavily on the manual crafting of structured responses. Future work could explore more automated methods for generating adversarial outputs, which would allow adversaries to exploit these vulnerabilities on a larger scale.
@@ 8 Conclusion
Appendix A Related work
LLM-based evaluation. Evaluating open-ended generation poses challenges due to the lack of a single valid ground truth. Human evaluation, though reliable, is expensive and time-consuming. To reduce costs and enable fast evaluation, powerful LLMs are often used as judges, LLM-based evaluators have been used for various specific tasks: providing AI feedback (Bai et al., 2022; Bubeck et al., 2023; Gudibande et al., 2023; Chiang et al., 2023; Zhou et al., 2023; Tan et al., 2023; Wang et al., 2023b; Kim et al., 2023; 2024; McAleese et al., 2024), evaluating text summarization (Gao et al., 2023; Luo et al., 2023), detecting LLM hallucination (Li et al., 2023a; Manakul et al., 2023; Adlakha et al., 2023; Cohen et al., 2023) etc. More recently, people have proposed to use powerful proprietary LLMs like GPT-4 to evaluate the general ability of LLMs as seen in benchmarks like G-eval (Liu et al., 2023b), MT-Bench and Chatbot Arena (Zheng et al., 2023), AlpacaEval (Dubois et al., 2023; Li et al., 2023c; Dubois et al., 2024), ArenaHard (Li et al., 2024c), WildBench (Lin et al., 2024), and MixEval (Ni et al., 2024).
Attacking LLM-based evaluations. While initially studied in the context of image classification, adversarial examples for language models have more recently been demonstrated for several tasks: question answering (Jia & Liang, 2017; Wallace et al., 2019), document classification (Ebrahimi et al., 2018), sentiment analysis (Alzantot et al., 2018; Maus et al., 2023), and toxicity (Jones et al., 2023; Wallace et al., 2019). More recently, Shi et al. (2023) found that LLM can be distracted by irrelevant context easily. Besides, there are also a lot of analyses to improve the robustness and reduce the bias of LLM-based evaluations. Liu et al. (2024) study the role of pairwise preferences in LLM evaluator alignment. Zheng et al. (2023) discusses the four limitations of LLM-as-a-Judge: position bias, verbosity bias, self-enhancement bias, and limited capability in grading math and reasoning questions. Regarding the verbosity bias, LLM judgers are known to be biased toward longer responses (Dubois et al., 2024; Zhao et al., 2024; Chen et al., 2024b).
Our work differs from these prior efforts in several aspects: 1) Unlike previous attacks that manipulate meaningful responses by appending adversarial suffixes, we propose the use of a completely non-informative “null model” that generates the same irrelevant output for all input instructions. This approach does not rely on producing contextually relevant responses, making it distinct from existing response-based adversarial attacks; 2) While many of the earlier works focus on optimizing individual prompts or attacks specific to a given input (Shi et al., 2024), our approach emphasizes the creation of universal, transferable adversarial prompts. These prompts are designed to work across various instructions without direct access to those instructions, offering a more generalized and powerful cheating strategy; 3) Most existing studies have focused on attacking open-source models or less-used benchmarks. To the best of our knowledge, no prior research has systematically targeted widely-used, state-of-the-art benchmarks like AlpacaEval 2.0 and Arena-Hard-Auto, or demonstrated the ability to achieve top-ranked win rates on these platforms. Our work presents the first comprehensive cheating on these highly influential LLM benchmarks.
Appendix B Implementation details